# A-VL: Adaptive Attention for Large Vision-Language Models

### v1

@@ Introduction
The decoder-only transformer-based LLM necessitates autoregressive iteration for output generation, where the generation of each token depends on all preceding tokens. This dependency significantly increases memory requirements and computational overhead. To address these challenges, existing methods are categorized into two main approaches: inference process optimization and computation reduction. Inference process optimization includes strategies such as page attention (Kwon et al. 2023) and speculative sampling (Xia et al. 2024) to achieve flexible and efficiency management of the entire inference process. Computation reduction involves methods like adaptive attention, early exit (Bae et al. 2023), and structured pruning (An et al. 2024). Our primary focus is on adaptive attention technology, which dynamically reduces computational redundancy according to inputs and has produced remarkable results in LLMs. StreamingLLM (Xiao et al. 2023) is designed to retain a fixed-position cache, which can train LLMs with long contexts on a limited cache window. Anagnostidis et al. (Anagnostidis et al. 2024) propose a new sigmoid function to dynamically sparse attention, but the model needs to be fine-tuned. FastGen (Ge et al. 2023) combines multiple strategies to retain recent caches, special characters, and caches that have been historically assigned heavier attention. H2(Zhang et al. 2024b) leverages historical accumulated attention and combined with the local cache for KV cache reduction, and has achieved remarkable results. Although these methods improve the inference efficiency of LLM, they are designed for single-modal text inputs. In addition, there are some studies focus on enhancing the efficiency of transformer-based visual models. SPViT (Kong et al. 2022) prunes tokens in Vision Transformers (ViT), and PuMer (Cao, Paranjape, and Hajishirzi 2023) prunes and merge tokens in cross attention; however, these methods are designed for encoder models. FastV (Chen et al. 2024) is an optimization solution designed for autoregressive LVLMs, revealing that without using KV cache, LVLM can potentially use only 50% of tokens after the second layer. Nevertheless, KV cache is an essential optimization tool for autoregressive inference. FastV has not been verified in inference with KV cache, because token pruning after the second layer may lead to the loss of some kv cache, making it impossible to select tokens in subsequent steps. And the pruning in FastV with KV cache may lead to the loss of potentially important tokens in the subsequent steps.
@@
where $A$ represents the attention weights assigned from the last token to all preceding tokens; $l$ denotes the index of the transformer decoder layers; and $h$ indicates the attention head. To investigate attention changes in LVLM, we randomly select 100 samples from the Flickr30k dataset (Young et al. 2014) that can output longer text sequences. We quantify the attention scores changes in LLaVA-1.5 7B model across three token types during the decode phase. The results are depicted in Figure 1. Surprisingly, the attention allocated to system input and image tokens gradually decreases as the token sequence lengthens, while the attention to user instructions decreases sharply. Note that user instructions are often of similar length to system inputs or even longer, but still exhibit different attention patterns. This suggests that the remote system inputs and image tokens are crucial for inference. These results indicate that the attention allocated to each modality may need to be managed separately.
@@
We use LLaVA-1.5 7B model to evaluate the OCRVQA dataset, and analyze the PPCI between the attention across each decoder layer and that in the first decoder layer. The results are depicted in Figure 4. As illustrated by the figure, the 50% PPCI between the first decoder layer and subsequent layers is less than 70%, and the 50% PPCI between the first layer and the last decoder layer is below 60%. We also calculate the 50% PPCI between the first three layers and the last layer, finding it to be less than 70% on average. Based on this observation, we refer to this phenomenon where different decoder layers focus on different image tokens as the heterogeneity of vision attention. Consequently, it is essential to design the adaptive attention method based on the attention distribution for each layers.
@@
We subsequently analyze the attention from the temporal dimension with the LLaVA-1.5 7B model and the OCRVQA dataset. During the autoregressive generation of tokens, we observe a high correlation in attention within the same decoder layer across different steps, as depicted in Figure 5. The figure illustrates that during the generation of 20 new tokens (20 steps) in the same decoder layer, the $p\%$ PPCI in image tokens between each step and first step. As depicted in the figure, each layer consistently allocates high attention to the top 50% of image tokens. For the top 30% image tokens, each layer allocates high attention in the short term, although this may shift after three steps. We refer to this short-term correlation of attention among image tokens within the same layer as the continuity of vision attention. This indicates that attention may gradually drift within partial image tokens. Given the gradual nature of these changes, we can reasonably assume that the set of core tokens remains stable in the short term.
@@
As illustrated in Figure 1, attention of text tokens decays rapidly. Therefore, it is advisable to focus on local text attention and a limited number of critical remote text attention. The H2O method can be precisely applied to achieve this objective. Specially, we maintain local text cache caches and evict useless text caches based on accumulated historical attention. We also analyze the attention patterns of Llama 2 (Touvron et al. 2023), a large language model. We observe that the attention patterns in Llama 2 are similar to those in the text token of LVLMs, both characterized by rapid attention decay. This similarity further underscores the suitability of employing an LLM’s adaptive attention method to the text tokens of LVLMs.
@@ A-VL Design
We design a hierarchical adaptive attention mechanism for vision attention based on the attention scores from the perspective of text tokens, as shown in Figure 6. In the prefill phase, we categorize image tokens into three sets: core, secondary, and minor, according to the attention score of the last text token. This classification uses the attention score from the text token to emphasize the semantic relationship between text and images. Specifically, after computing the attention score in each decoder layer, tokens within the top $C\%$ of scores are designated as core tokens, those within the top $S\%$ as secondary tokens, and the remainder as minor tokens. We ensure that $S>C$, thereby making core tokens a subset of secondary tokens. Minor tokens are immediately evicted to minimize KV cache usage. Secondary tokens, which may be useful in future, are retained in memory; however, they are not all used in computations. Only core tokens are involved in the computations of subsequent attention modules. However, our observations suggest that vision attention may shift gradually among secondary tokens, potentially altering the set of core tokens. Thus, every $K$ steps, we utilize all secondary tokens for inference and compute the attention scores to update the core token set, ensuring that attention remains focused on the most significant tokens. We refer to this step as the update decode phase. This design is based on our insights into the continuity of vision attention, allowing the model to minimize cache usage and further reduce computational load of attention. It should be noted that our method is applied independently to each decoder layer, and the results are not utilized until running to the corresponding layer in the next step, allowing for parallel processing with the original model’s inference.
@@
Adaptive Text Attention
Based on the insight of text attention, when generating new text tokens, both remote image tokens and the local text tokens are used to predict. Since image tokens are processed separately, we use H2O’s accumulated attention method for text cache eviction. Specially, we set a text cache window size of $T\%$ of the maximum text token length. After computing attention weights at each layer, we separate the attention of vision and text, and store the accumulated text attention scores. If the text cache size exceed the set text cache window size, we evict the redundant caches from the first half of the window based on these scores. This method, utilized by H2O, has proven effective in LLM. We incorporate this approach into the text attention of LVLMs and verify its effectiveness through experiments. We list all the parameters involved in our method and their description in Table 1.
Figure 7: Adaptive vision attention use only core KV cache. The details such as softmax are omitted in the figure.
Model Method DocVQA TextVQA OCRBench Nocaps Flickr30k Avg.
LLaVA 1.5 7B Original 31.20 48.67 20.60 101.88 73.81 55.23
FastV 27.35 47.16 19.30 101.06 72.99 53.57
H2O 30.34 48.22 20.40 102.23 73.53 54.94
Our 31.26 48.44 20.90 102.67 73.81 55.42
LLaVA 1.5 13B Original 35.39 52.86 22.50 107.53 79.69 59.59
FastV 30.27 50.64 20.60 106.41 78.88 57.36
H2O 33.86 52.72 22.30 107.11 79.15 59.03
Our 35.22 52.97 22.40 107.52 79.79 59.58
LLaVA 1.6 7B Original 74.43 64.65 53.10 88.18 68.45 69.76
FastV 67.69 63.15 48.30 86.21 66.73 66.42
H2O 72.01 64.15 51.10 88.63 68.61 68.90
Our 74.46 64.69 52.00 88.14 68.38 69.53
LLaVA 1.6 13B Original 77.23 67.06 53.90 88.12 66.67 70.60
FastV 70.26 65.14 49.10 87.78 65.96 67.64
H2O 75.85 66.58 53.00 88.33 66.71 70.09
Our 77.30 66.97 53.90 87.68 66.52 70.48
Table 2: Performance comparison of different models, different tasks and different methods.
@@
Model Settings Stored Cache Used Cache DocVQA TextVQA OCRBench Nocaps
@@
LLaVA 1.5 7B - - - - - 100% 100% 31.20 48.67 20.60 101.88
40 - - - - 46% 46% 31.26 48.44 20.90 102.67
40 30 3 - - 46% 39% 31.02 48.44 21.00 102.67
45 30 3 90 - 50% 40% 31.02 48.46 20.70 102.34
45 30 3 90 50 48% 38% 30.92 48.46 20.50 102.22
LLaVA 1.5 13B - - - - - 100% 100% 35.39 52.86 22.50 107.53
40 - - - - 46% 46% 35.37 52.86 22.20 107.61
45 30 3 - - 50% 40% 35.22 52.97 22.40 107.52
45 30 3 90 - 50% 40% 35.46 52.83 22.30 107.60
45 30 3 90 60 48% 38% 35.43 52.84 22.30 107.59
LLaVA 1.6 7B - - - - - 100% 100% 74.43 64.65 53.10 88.18
40 - - - - 43% 43% 74.29 64.76 51.20 88.14
45 30 3 90 - 47% 36% 74.46 64.71 51.90 87.43
45 30 3 90 70 45% 35% 74.44 64.69 52.00 87.45
LLaVA 1.6 13B - - - - - 100% 100% 77.23 67.06 53.90 88.12
40 - - - - 43% 43% 77.30 66.97 53.90 87.68
45 30 3 90 - 47% 36% 76.97 67.06 54.40 87.59
45 30 3 90 70 45% 35% 76.97 67.06 54.50 87.59
Table 3: Performance under different settings with our method.
@@
We employ multiple vision-language tasks to evaluate our method. Since our method primarily optimizes the KV cache in the decode phase, we do not choice any multiple-choice datasets that generate only one token during the prefill phase. The tasks and datasets used in our evaluation are as follows:
@@
  • Visual Question Answering (VQA). This task requires models to answer questions based on visual information from images. We use two datasets, DocVQA (Mathew, Karatzas, and Jawahar 2021) and TextVQA (Singh et al. 2019), measuring performance with the ANLS (Biten et al. 2019) and accuracy metrics respectively.
@@
We use mainstream LVLMs and their different sizes, including LLaVA 1.5 7B, LLaVA 1.5 13B, LLaVA 1.6 7B and LLaVA 1.6 13B.
@@
We compare the output of the original model, FastV designed for LVLMs, and H2O designed for LLMs. For the H2O method, set the cache window to 75% recent and 25% heavy hitter. For FastV, in order to compare performance with KV cache, we delete redundant tokens according to the FastV method in the prefill phase, and use the generated KV cache for inference in the decode phase.
Performance Guarantee
We evaluate the performance of various models across different tasks, as shown in Table 2. FastV, H2O, and our method retains 50% of the KV cache. Our method achieves nearly lossless performance using only 50% of the KV cache. And our average performance is better than the other two baselines. FastV shows reduced performance when utilizing the KV cache, due to the loss of potentially useful tokens during token pruning. The H2O method designed for LLMs shows effectiveness, but its average performance is still inferior to our method. Our method demonstrates good results across various models, confirming its effectiveness.
@@
Figure 9: Performance under different parameter S, C, K.
@@
We first analyze the relationship between KV cache memory usage and performance across various parameters with LLaVA-1.5 7B model and DocVQA dataset. For vision attention, we evaluate the performance with different parameters, as shown in Figure 9. Figure 9(a) illustrates that retaining 40% of the image cache as the secondary cache guarantees no performance loss, and retaining only 20% results can ensure performance decline within 3%. According to Figure 9(b), computing with only the core cache yields better results than simply eliminating these caches. Additionally, increasing the update frequency (lowering K) ensures that the core cache consistently contains the most critical elements. These results confirms the insights previously discussed. Performance under varying parameters P and T are presented in Figure 10. Specifically, maintaining over 60% of the text cache (Parameter T) ensures stable performance. In contrast, retaining less than 80% of the image tokens during the prefill phase (Parameter P) leads to performance degradation. This underscores the underperformance of the FastV method with the KV cache, as previously analyzed. However, integration with our method can reduce computational load by nearly 10% during the prefill phase.
Figure 10: Performance under different parameter P and T.

### v2

@@ Introduction
The decoder-only transformer-based LLM necessitates autoregressive iteration for output generation, where the generation of each token depends on all preceding tokens. This dependency significantly increases memory requirements and computational overhead. To address these challenges, existing methods are categorized into two main approaches: inference process optimization and computation reduction. Inference process optimization includes strategies such as page attention (Kwon et al. 2023) and speculative sampling (Xia et al. 2024) to achieve flexible and efficient management of the entire inference process. Computation reduction involves methods like adaptive attention, early exit (Bae et al. 2023), and structured pruning (An et al. 2024). Our primary focus is on adaptive attention technology, which dynamically reduces computational redundancy according to inputs and has produced remarkable results in LLMs. StreamingLLM (Xiao et al. 2023) is designed to retain a fixed-position cache, which can train LLMs with long contexts on a limited cache window. Anagnostidis et al. (Anagnostidis et al. 2024) propose a new sigmoid function to dynamically sparse attention, but the model needs to be fine-tuned. FastGen (Ge et al. 2023) combines multiple strategies to retain recent caches, special characters, and caches that have been historically assigned heavier attention. H2(Zhang et al. 2024b) leverages historical accumulated attention and combined with the local cache for KV cache reduction, and has achieved remarkable results. Although these methods improve the inference efficiency of LLM, they are designed for single-modal text inputs. In addition, there are some studies that focus on enhancing the efficiency of transformer-based visual models. SPViT (Kong et al. 2022) prunes tokens in Vision Transformers (ViT), and PuMer (Cao, Paranjape, and Hajishirzi 2023) prunes and merge tokens in cross attention; however, these methods are designed for encoder models. FastV (Chen et al. 2024) is an optimization solution designed for autoregressive LVLMs, revealing that without using KV cache, LVLM can potentially use only 50% of tokens after the second layer. Nevertheless, the KV cache is an essential optimization tool for autoregressive inference. FastV has not been verified in inference with KV cache, because token pruning after the second layer may lead to the loss of some KV cache, making it impossible to select tokens in subsequent steps. And the pruning in FastV with the KV cache may lead to the loss of potentially important tokens in the subsequent steps.
@@
where $A$ represents the attention weights assigned from the last token to all preceding tokens; $l$ denotes the index of the transformer decoder layers; and $h$ indicates the attention head. To investigate attention changes in LVLM, we randomly select 100 samples from the Flickr30k dataset (Young et al. 2014) that can output longer text sequences. We quantify the attention score changes in LLaVA-1.5 7B model across three token types during the decode phase. The results are depicted in Figure 1. Surprisingly, the attention allocated to system input and image tokens gradually decreases as the token sequence lengthens, while the attention to user instructions decreases sharply. Note that user instructions are often of similar length to system inputs or even longer, but still exhibit different attention patterns. This suggests that the remote system inputs and image tokens are crucial for inference. These results indicate that the attention allocated to each modality may need to be managed separately.
@@
We use LLaVA-1.5 7B model to evaluate the OCRVQA dataset, and analyze the PPCI between the attention across each decoder layer and that in the first decoder layer. The results are depicted in Figure 4. As illustrated by the figure, the 50% PPCI between the first decoder layer and subsequent layers is less than 70%, and the 50% PPCI between the first layer and the last decoder layer is below 60%. We also calculate the 50% PPCI between the first three layers and the last layer, finding it to be less than 70% on average. Based on this observation, we refer to this phenomenon where different decoder layers focus on different image tokens as the heterogeneity of vision attention. Consequently, it is essential to design the adaptive attention method based on the attention distribution for each layer.
@@
We subsequently analyze the attention from the temporal dimension with the LLaVA-1.5 7B model and the OCRVQA dataset. During the autoregressive generation of tokens, we observe a high correlation in attention within the same decoder layer across different steps, as depicted in Figure 5. The figure illustrates that during the generation of 20 new tokens (20 steps) in the same decoder layer, the $p\%$ PPCI in image tokens between each step and the first step. As depicted in the figure, each layer consistently allocates high attention to the top 50% of image tokens. For the top 30% image tokens, each layer allocates high attention in the short term, although this may shift after three steps. We refer to this short-term correlation of attention among image tokens within the same layer as the continuity of vision attention. This indicates that attention may gradually drift within partial image tokens. Given the gradual nature of these changes, we can reasonably assume that the set of core tokens remains stable in the short term.
@@
As illustrated in Figure 1, the attention of text tokens decays rapidly. Therefore, it is advisable to focus on local text attention and a limited number of critical remote text attention. The H2O method can be precisely applied to achieve this objective. Specially, we maintain local text cache caches and evict useless text caches based on accumulated historical attention. We also analyze the attention patterns of Llama 2 (Touvron et al. 2023), a large language model. We observe that the attention patterns in Llama 2 are similar to those in the text token of LVLMs, both characterized by rapid attention decay. This similarity further underscores the suitability of employing an LLM’s adaptive attention method to the text tokens of LVLMs.
@@ A-VL Design
We design a hierarchical adaptive attention mechanism for vision attention based on the attention scores from the perspective of text tokens, as shown in Figure 6. In the prefill phase, we categorize image tokens into three sets: core, secondary, and minor, according to the attention score of the last text token. This classification uses the attention score from the text token to emphasize the semantic relationship between text and images. Specifically, after computing the attention score in each decoder layer, tokens within the top $C\%$ of scores are designated as core tokens, those within the top $S\%$ as secondary tokens, and the remainder as minor tokens. We ensure that $S>C$, thereby making core tokens a subset of secondary tokens. Minor tokens are immediately evicted to minimize KV cache usage. Secondary tokens, which may be useful in the future, are retained in memory; however, they are not all used in computations. Only core tokens are involved in the computations of subsequent attention modules. However, our observations suggest that vision attention may shift gradually among secondary tokens, potentially altering the set of core tokens. Thus, every $K$ steps, we utilize all secondary tokens for inference and compute the attention scores to update the core token set, ensuring that attention remains focused on the most significant tokens. We refer to this step as the update decode phase. This design is based on our insights into the continuity of vision attention, allowing the model to minimize cache usage and further reduce computational load of attention. It should be noted that our method is applied independently to each decoder layer, and the results are not utilized until running to the corresponding layer in the next step, allowing for parallel processing with the original model’s inference.
@@
Adaptive Text Attention
Based on the insight of text attention, when generating new text tokens, both remote image tokens and local text tokens are used to predict. Since image tokens are processed separately, we use H2O’s accumulated attention method for text cache eviction. Specially, we set a text cache window size of $T\%$ of the maximum text token length. After computing attention weights at each layer, we separate the attention of vision and text, and store the accumulated text attention scores. If the text cache size exceeds the set text cache window size, we evict the redundant caches from the first half of the window based on these scores. This method, utilized by H2O, has proven effective in LLM. We incorporate this approach into the text attention of LVLMs and verify its effectiveness through experiments. We list all the parameters involved in our method and their description in Table 1.
Figure 7: Adaptive vision attention uses only core KV cache. The details such as softmax are omitted in the figure.
@@
Model Method DocVQA TextVQA OCRBench Nocaps Flickr30k VQAv2 Avg.
LLaVA 1.5 7B Original 31.20 48.67 20.60 101.88 73.81 76.52 58.78
FastV 27.35 47.16 19.30 101.06 72.99 74.86 57.12
H2O 30.34 48.22 20.40 102.23 73.53 76.49 58.54
Our 31.26 48.44 20.90 102.67 73.81 76.52 58.93
LLaVA 1.5 13B Original 35.39 52.86 22.50 107.53 79.69 77.97 62.66
FastV 30.27 50.64 20.60 106.41 78.88 76.65 60.57
H2O 33.86 52.72 22.30 107.11 79.15 77.96 62.18
Our 35.22 52.97 22.40 107.52 79.79 77.97 62.65
LLaVA 1.6 7B Original 74.43 64.65 53.10 88.18 68.45 80.01 71.47
FastV 67.69 63.15 48.30 86.21 66.73 79.58 68.61
H2O 72.01 64.15 51.10 88.63 68.61 79.97 70.75
Our 74.46 64.69 52.00 88.14 68.38 80.02 71.28
LLaVA 1.6 13B Original 77.23 67.06 53.90 88.12 66.67 80.93 72.32
FastV 70.26 65.14 49.10 87.78 65.96 80.48 69.79
H2O 75.85 66.58 53.00 88.33 66.71 80.91 71.90
Our 77.30 66.97 53.90 87.68 66.52 80.94 72.22
Qwen-VL Original 68.65 61.10 49.10 58.92 59.01 79.17 62.66
FastV 51.45 52.85 38.60 58.73 57.84 76.05 55.92
H2O 65.84 60.59 46.60 59.06 59.39 79.16 61.77
Our 68.09 60.66 49.10 58.46 58.88 79.16 62.39
Table 2: Performance comparison of different models, different tasks and different methods. Except for the original model, each method only retains 50% of the KV cache. The first and second highest average metrics are bolded.
@@
Model Settings Stored Cache Used Cache KV Cache Memory DocVQA TextVQA OCRBench Nocaps
@@
LLaVA 1.5 7B - - - - - 100% 100% 339 MB 31.20 48.67 20.60 101.88
40 - - - - 46% 46% 156 MB 31.26 48.44 20.90 102.67
40 30 3 - - 46% 39% 156 MB 31.02 48.44 21.00 102.67
45 30 3 90 - 50% 40% 170 MB 31.02 48.46 20.70 102.34
45 30 3 90 50 48% 38% 163 MB 30.92 48.46 20.50 102.22
LLaVA 1.5 13B - - - - - 100% 100% 529 MB 35.39 52.86 22.50 107.53
40 - - - - 46% 46% 243 MB 35.37 52.86 22.20 107.61
45 30 3 - - 50% 40% 265 MB 35.22 52.97 22.40 107.52
45 30 3 90 - 50% 40% 265 MB 35.46 52.83 22.30 107.60
45 30 3 90 60 48% 38% 254 MB 35.43 52.84 22.30 107.59
LLaVA 1.6 7B - - - - - 100% 100% 1179 MB 74.43 64.65 53.10 88.18
40 - - - - 43% 43% 507 MB 74.29 64.76 51.20 88.14
45 30 3 90 - 47% 36% 554 MB 74.46 64.71 51.90 87.43
45 30 3 90 70 45% 35% 531 MB 74.44 64.69 52.00 87.45
LLaVA 1.6 13B - - - - - 100% 100% 1841 MB 77.23 67.06 53.90 88.12
40 - - - - 43% 43% 792 MB 77.30 66.97 53.90 87.68
45 30 3 90 - 47% 36% 865 MB 76.97 67.06 54.40 87.59
45 30 3 90 70 45% 35% 828 MB 76.97 67.06 54.50 87.59
Table 3: Performance under different settings with our method. KV Cache Memory refers to the memory usage when the batch size is one. As the batch size increases, memory consumption scales proportionally, resulting in greater memory savings.
@@
We employ multiple vision-language tasks to evaluate our method. Since our method primarily optimizes the KV cache in the decode phase, we do not choose any multiple-choice datasets that generate only one token during the prefill phase. For example, SEED-Bench (Li et al. 2024) is a fully multiple-choice benchmark, so we did not select it. The tasks and datasets used in our evaluation are as follows:
@@
  • Visual Question Answering (VQA). This task requires models to answer questions based on visual information from images. We use three datasets, DocVQA (Mathew, Karatzas, and Jawahar 2021), TextVQA (Singh et al. 2019) and VQAv2 (Goyal et al. 2017), measuring performance with the ANLS (Biten et al. 2019) and accuracy metrics.
@@
Among these tasks, some emphasize the overall content of the image (e.g., image caption), and some focus on fine-grained details (e.g., OCR). This distinction enables the evaluation across varying levels of granularity.
@@
We use mainstream LVLMs and their different sizes, including LLaVA 1.5 7B, LLaVA 1.5 13B, LLaVA 1.6 7B, LLaVA 1.6 13B and Qwen-VL.
@@
We compare the output of the original model, FastV designed for LVLMs, and H2O designed for LLMs. For the H2O method, set the cache window to 75% recent and 25% heavy hitter. For FastV, in order to compare performance with the KV cache, we delete redundant tokens according to the FastV method in the prefill phase, and use the generated KV cache for inference in the decode phase.
@@
Figure 9: Performance under different parameters S, C, K.
Performance Guarantee
We evaluate the performance of various models across different tasks, as shown in Table 2. FastV, H2O, and our method retains 50% of the KV cache. Our method achieves nearly lossless performance using only 50% of the KV cache. And our average performance is better than the other two baselines. FastV shows reduced performance when utilizing the KV cache, due to the loss of potentially useful tokens during token pruning. The H2O method designed for LLMs shows effectiveness, but its average performance is still inferior to our method. We observe that while H2O can sometimes maintain performance like our method in image caption tasks that focus on the overall image information, our method demonstrates better performance in OCR and VQA tasks, which require attention to finer details. This indicates that reducing the number of image tokens may have minimal impact on tasks emphasizing general image content, but for detail-oriented tasks, our method, specifically designed for LVLMs is essential. Our method demonstrates good results across various models, confirming its effectiveness. In all methods, 50% of the KV Cache is retained, resulting in similar decoding inference latency. Detailed memory savings from the KV Cache are shown in Table 3.
Figure 10: Performance under different parameters P and T.
@@
We first analyze the relationship between KV cache memory usage and performance across various parameters with LLaVA-1.5 7B model and DocVQA dataset. For vision attention, we evaluate the performance with different parameters, as shown in Figure 9. Figure 9(a) illustrates that retaining 40% of the image cache as the secondary cache guarantees no performance loss, and retaining only 20% results can ensure performance decline within 3%. According to Figure 9(b), computing with only the core cache yields better results than simply eliminating these caches. Additionally, increasing the update frequency (lowering K) ensures that the core cache consistently contains the most critical elements. These results confirm the insights previously discussed. Performance under varying parameters P and T are presented in Figure 10. Specifically, maintaining over 60% of the text cache (Parameter T) ensures stable performance. In contrast, retaining less than 80% of the image tokens during the prefill phase (Parameter P) leads to performance degradation. This underscores the underperformance of the FastV method with the KV cache, as previously analyzed. However, integration with our method can reduce computational load by nearly 10% during the prefill phase.